a.) Supposed that the length of the pregnancy is normally distribute with the mean of 266 days and the standard deviation of 14 days. What is the probability that a randomly chosen pregnancy will last at least 258 days

b.)Supposed that the length of the pregnancy is normally distribute with the mean of 266 days and the standard deviation of 14 days. What is the probability that a randomly chosen pregnancy will last fewer than 283 days? Round your answer to the three digits.
Answer #1

a) P(Pregnancy will last at least 258 days)

= P(z > (258 - 266)/14)

= P(z > -0.57)

= 0.716

b) P(Pregnancy will last fewer than 283 days)

= P(z < (283 - 266)/14)

= P(z < 1.21)

= 0.888